# Emergent Hierarchical Reasoning in LLMs through Reinforcement Learning

### v1

@@ 1 Introduction
This discovery – that the learning frontiers dynamically shifts to strategic planning – is more than an academic curiosity; it provides a clear blueprint for a more effective RL algorithm. If the primary driver for advanced reasoning is the mastery of high-level strategic planning, then current agnostic credit assignment methods used in the prevailing GRPO (Guo et al., 2025) and its variants (Yu et al., 2025; Liu et al., 2025b; Wang et al., 2025b) are fundamentally inefficient, as they dilute optimization pressure across all tokens rather than concentrating it where it matters most.
@@ 2.2.1 Forging a Reliable Procedural Engine
We indeed find that the phase of procedural consolidation might not appear in models with stronger capacity, as evident in MiMO-VL-Instruct and Qwen-4B-Instruct (see Appendix). This leads us to the belief that the primary driver of RL is indeed the exploration of strategic planning.
@@ 2.2.2 Steering the Skills with Strategic Planning
The third column of Figure 2 provides clear evidence of this strategic exploration phase. The semantic entropy of strategic grams (red line, marked with ②) shows a distinct and steady increase. This indicates that the model is not converging on a single optimal strategy but is instead actively expanding its repertoire of strategic plans. This observation is critical: mastery in reasoning, in this context, is achieved by developing a rich and varied strategic playbook, which contrasts sharply with the collapse in token-level entropy seen during the initial procedural consolidation phase.
@@
Once payoff from procedural consolidation diminishes, performance gains are driven primarily by exploring high-level strategies. This is reflected in the increasing semantic diversity of strategic grams, which correlate with reasoning performance and represents the key learning frontier.
@@
  • ”Aha moments” are the behavioral signature of the model discovering, mastering, and reinforcing a new, powerful strategy or set of strategic constructs.
  • ”Length-scaling” is highly consistent with increase in strategic diversity. As Figure 2 shows, the rise in semantic entropy of planning tokens is strongly correlated with an increase in average sequence length. More sophisticated strategies – involving planning, case analysis, and self-reflections – are mediated by planning tokens and naturally produce longer, more successful reasoning traces.
@@
Unluckily, the decrease in token-level entropy sometimes mislead practitioners into the conception of declined exploration. This is incorrect, however, as it contradicts the fact of increasing exploration in strategic plans (semantic entropy of planning tokens) and the improving reasoning performance.
@@ 3 HICRA: Hierarchy-Aware Credit Assignment
The core mechanism of HICRA engineers more effective exploration by reshaping the policy update’s target distribution. A standard policy gradient (Williams, 2004) update nudges the policy $\pi_{\theta_{old}}$ toward an implicit target distribution $\pi^{\ast}$ defined by the advantage function:
$\pi^{\ast}(o_{i,t}|\mathbf{q},\mathbf{o}_{i,<t})\propto\pi_{\theta_{old}}(o_{i,t}|\mathbf{q},\mathbf{o}_{i,<t})\exp(\eta\hat{A}_{i,t})$
where $\eta$ is the learning rate. Typically, this update pressure is applied isotropically, affecting all token types uniformly. HICRA breaks this symmetry. By using the modified advantage $\hat{A}^{\text{HICRA}}$, it creates a new target distribution, $\pi^{*}_{\text{HICRA}}$, that is anisotropically stretched toward the strategic dimensions of the action space. This new target distribution places significantly greater probability mass on planning tokens, particularly those within high-reward trajectories.
@@ 4.1 Experimental Setup
Our experiments are conducted on four open-source model families, including Qwen2.5-7B (Yang et al., 2024), Qwen3-4B (Yang et al., 2025), LLama-3.1-8B (Grattafiori et al., 2024), and MiMO-VL-7B (Xiaomi, 2025). We use both base and instruction-tuned variants to ensure our findings are consistent and test generalization to vision-language models.
Benchmarks
We evaluate all models on a diverse set of challenging mathematical reasoning benchmarks to rigorously test their complex reasoning capabilities. The benchmarks include AIME24, AIME25 (Mathematical Association of America, 2024), Math500 (Lightman et al., 2023), AMC23, Minerva (Lewkowycz et al., 2022), and Olympiad (He et al., 2024). We follow Deepseek R1 (Guo et al., 2025)’s evaluation protocol, where the performance is measured by Pass@1 Accuracy with temperature $0.6$ sampling.
@@
We compare HICRA against three primary baselines: the Base model (before RL), the widely adopted GRPO baseline with clip-higher (Yu et al., 2025), and Entropy Regularization: GRPO with an additional regularization loss on token-level entropy (Cheng et al., 2025). For HICRA, we set the amplification hyperparameter $\alpha$ to $0.2$ and identify planning tokens using the Strategic Grams (SGs) methodology detailed in Section 2.1.
@@ 4.2 Main Results
QW3-4B-INSTRUCT (NO-THINK)
@@
QW2.5-7B-BASE
@@
     ORZ (Hu et al., 2025) 18.8 14.8 80.2 55.1 38.6 45.9
     SimpleRL (Zeng et al., 2025) 16.7 3.3 78.2 40.4
     Entropy Regularization 16.0 9.3 57.4 54.3
@@
Our primary results, summarized in Table 1, show that HICRA consistently and significantly outperforms both the Base and GRPO baselines across all models and benchmarks. On the strongest base model, Qwen3-4B-Instruct, HICRA’s gains demonstrate that even on highly capable models, selectively amplifying the learning signal for strategic reasoning yields substantial improvements. This trend holds for non-instruct-tuned models as well, providing strong empirical evidence for our central claim: by identifying and focusing on the emergent strategic bottleneck, HICRA accelerates the development of advanced reasoning abilities more efficiently than agnostic methods.
@@ 4.3.3 HICRA’s Limitation: Dependency on a Procedural Foundation
HICRA’s effectiveness is predicated on a key assumption: that the base model should readily possess a reasonable foundation for low-level procedural correctness. As shown in Figure 6, when this foundation is lacking – as observed with Llama-3.1-Instruct – HICRA can fail to provide an advantage over GRPO. Seen from the semantic entropy graph in, there is a reverse trend between GRPO and HICRA, implying an opposite training focus on planning tokens and execution tokens. HICRA’s enforced strategic exploration becomes counterproductive if the model cannot reliably execute the plans it generates, leading to unstable learning dynamics and learning effects observed on Llama-3.1. This suggests that HICRA is most effective when applied to models that have already achieved a degree of procedural reliability, highlighting an important dependency for future work on more adaptive, model-aware hierarchical methods.
4.4 A Compass for Strategic Exploration
@@
As shown in Figure 7 for MiMO-VL-7B, token-level entropy collapses for both HICRA and GRPO, which might naively suggest that exploration has ceased. Similarly, the Pass@8 (Training) metric quickly saturates, rendering it useless for distinguishing the ongoing learning dynamics.
@@
Recent work has proposed high-entropy tokens, sometimes called “fork tokens,” to imply its role as proxies for decision points in a reasoning trace (Wang et al., 2025c). Our analysis investigates the relationship between our functionally-defined planning tokens and this entropy-based definition.
@@ 5 Related Works
Reinforcement Learning for LLM Reasoning. The application of Reinforcement Learning (RL) has been pivotal in enhancing the complex reasoning abilities of Large Language Models (LLMs). Seminal work by Ouyang et al. demonstrated the effectiveness of learning from human feedback to align models with user instructions. More recently, algorithms like Group Reward Policy Optimization (Guo et al., 2025) have been developed to specifically incentivize reasoning capabilities in LLMs (Liu et al., 2025b; Yu et al., 2025; Team et al., 2025; Liu et al., 2025a), VLMs (Wang et al., 2025b), Tool-Use Agents (Wang et al., 2025a; Su et al., 2025), leading to significant performance gains on downstream performance. While these methods have proven empirically successful, they typically apply optimization pressure agnostically across all generated tokens, without distinguishing between different functional roles within the reasoning process. Our work builds on this foundation but introduces a more targeted approach by focusing on the emergent reasoning hierarchy.
@@
Furthermore, recent work has identified high-entropy ”fork tokens” as potential proxies for critical decision points in reasoning (Wang et al., 2025c). Our work distinguishes itself by defining planning tokens based on their semantic function. We also validate the limitation of identifying crucial tokens solely based on entropy.
@@
Hierarchical Reasoning and Cognition. The concept of hierarchical processing is a cornerstone of cognitive neuroscience, which posits that the human brain separates high-level, abstract planning from low-level motor or procedural execution (Huntenburg et al., 2018; Murray et al., 2014; Zeraati et al., 2023). This cognitive architecture provides a compelling parallel to the functional hierarchy we identify in RL-tuned LLMs. Our research is inspired by this parallel, proposing that LLMs similarly develop a functional separation between strategic planning and procedural execution.
@@ 6 Conclusion and Promising Directions
Our findings open several promising avenues for future research:
Finer-Grained Credit Assignment: Our analyses and results highlight the value of fine-grained credit assignment, e.g., process reward models, especially for high-impact tokens.
Hierarchical Methods: The failure of HICRA on Llama-3.1-Instruct implies the mutual roles of procedural reliability and strategic planning. It is worthwhile to study how to effectively facilitate both roles and adaptively diagnoses the learning bottleneck across different models.
@@
Generalizing the Reasoning Hierarchy: While our analysis focused on mathematical reasoning, the concept of a functional hierarchy between planning and execution is likely a general feature of complex, multi-step tasks. Investigating whether similar dynamics emerge in domains like code generation, scientific discovery, and long-form text planning could yield novel, domain-specific optimization strategies.

### v2

@@ Abstract
Figure 1: (Left) LLM reasoning mirrors a human-like hierarchical reasoning: high-level strategic planning and low-level procedural executions. (Right) Hierarchical reasoning emerges during RL training via a two-phase dynamic. Phase ① consolidates low-level skills, marked by a token-entropy drop in execution tokens. The learning frontier then shifts to Phase ②, where the model explores and masters high-level planning, marked by increased semantic diversity, sustained reasoning enhancement and length scaling.
@@ 1 Introduction
This discovery – that the learning frontiers dynamically shifts to strategic planning – is more than an academic curiosity; it provides a clear blueprint for a more effective RL algorithm. If the primary driver for advanced reasoning is the mastery of high-level strategic planning, then current agnostic credit assignment methods used in the prevailing GRPO (Guo et al., 2025) and its variants (Yu et al., 2025; Liu et al., 2025b; Wang et al., 2025c) are fundamentally inefficient, as they dilute optimization pressure across all tokens rather than concentrating it where it matters most.
@@ 2.2.1 Forging a Reliable Procedural Engine
We indeed find that the phase of procedural consolidation might be absent or shot in models with stronger capacity, as evident in MiMO-VL-Instruct and Qwen-4B-Instruct. We refer the reader to check the full analysis of training dynamics across six models in the appendix. This leads us to the belief that the primary driver of RL is indeed the exploration of strategic planning.
@@ 2.2.2 Steering the Skills with Strategic Planning
The third column of Figure 3 provides clear evidence of this strategic exploration phase. The semantic entropy of strategic grams (red line, marked with ②) shows a distinct and steady increase. This indicates that the model is not converging on a single optimal strategy but is instead actively expanding its repertoire of strategic plans. This observation is critical: mastery in reasoning, in this context, is achieved by developing a rich and varied strategic playbook, which contrasts sharply with the sharp decrease in token-level entropy seen during the initial procedural consolidation phase.
@@
Once payoff from procedural consolidation diminishes, performance gains are driven primarily by exploring high-level strategies. This is marked by the increasing semantic diversity of strategic grams, which correlate with sustained reasoning enhancement, length scaling, and represents the key learning frontier.
@@
  • “Aha moments” are the behavioral signature of the model discovering, mastering, and reinforcing a new, powerful strategy or set of strategic constructs.
  • “Length-scaling” is highly consistent with increase in strategic diversity. As Figure 3 shows, the rise in semantic entropy of planning tokens is strongly correlated with an increase in average sequence length. More sophisticated strategies – involving planning, case analysis, and self-reflections – are mediated by planning tokens and naturally produce longer, more successful reasoning traces.
@@
Unluckily, the decrease in token-level entropy sometimes mislead practitioners into the conception of declined exploration. This is incorrect, however, as it contradicts the fact of increasing exploration in strategic plans (semantic entropy of planning tokens) and the improving reasoning performance. We refer interested readers to the appendix for a full analysis of training dynamics across six LLM and VLMs in the appendix.
Semantic entropy avoids this pitfall by directly measuring diversity at the semantic level of meaningful strategic units. Its trend accurately reflects the expansion of the model’s strategic playbook, making it a more reliable diagnostic tool for tracking genuine exploration and predicting sustained performance improvements. In Section 4.4, we showcase how semantic entropy avoids the flaws in token entropy and complements Pass@K metric with further benefits.
@@ 3 HICRA: Hierarchy-Aware Credit Assignment
The core mechanism of HICRA engineers more effective exploration by reshaping the policy update’s target distribution. A standard policy gradient (Williams, 2004) update nudges the policy $\pi_{\theta_{old}}$ toward an implicit target distribution $\pi^{\ast}$ defined by the advantage function described as follows (the derivation is included in the appendix):
$\pi^{\ast}(o_{i,t}|\mathbf{q},\mathbf{o}_{i,<t})\propto\pi_{\theta_{old}}(o_{i,t}|\mathbf{q},\mathbf{o}_{i,<t})\exp(\hat{A}_{i,t})$
Typically, this update pressure is applied isotropically, affecting all token types uniformly. HICRA breaks this symmetry. By using the modified advantage $\hat{A}^{\text{HICRA}}$, it creates a new target distribution, $\pi^{*}_{\text{HICRA}}$, that is anisotropically stretched toward the strategic dimensions of the action space. This new target distribution places significantly greater probability mass on planning tokens (through the term $\text{exp}(\hat{A}_{i,t})$), particularly those within high-reward trajectories.
@@ 4.1 Experimental Setup
Our experiments are conducted on four open-source model families, including Qwen2.5-7B (Yang et al., 2024), Qwen3-4B (Yang et al., 2025), LLama-3.1-8B (Grattafiori et al., 2024), Qwen2.5-VL-7b (Yang et al., 2024) and MiMO-VL-7B (Xiaomi, 2025). We use both base and instruction-tuned variants to ensure our findings are consistent and test generalization to vision-language models.
Training Datasets and Benchmarks
[…]
@@
We compare HICRA against three primary baselines: the Base model (before RL), the widely adopted GRPO baseline with clip-higher (Yu et al., 2025) by default, and Entropy Regularization: GRPO with an additional regularization loss on token-level entropy (Cheng et al., 2025). For HICRA, we set the amplification hyperparameter $\alpha$ to $0.2$ and identify planning tokens using the Strategic Grams (SGs) methodology detailed in Section 2.1. For all experiments, we increase the training context length from 16K to 32K when the response clip rate exceeds 20% (Luo et al., 2025). For the specific experiments on Llama-3.1-Instruct, we add a dynamic filtering mechanism (Yu et al., 2025) based on GRPO Clip-Higher due to significant vanishing advantanges (Wang et al., 2025c).
@@ 4.2 Main Results
Qwen3-4B-Instruct-2507
     Base 63.4 47.7 94.6 86.7 45.2 72.4
     GRPO 68.5 60.0 96.2 88.5 50.0 72.7
     HICRA 73.1 65.1 97.2 90.2 50.7 72.0
     $\Delta$ (HICRA - GRPO) +5.4 +5.1 +1.0 +1.7 +0.7 -0.7
Qwen3-4B-Adaptive (No-Think)
@@
Llama-3.1-8B-Instruct
     Base 4.2 0.6 50.2 17.1 20.9 13.7
     GRPO 8.9 0.5 53.0 25.0 27.2 20.3
     HICRA 8.3 0.8 54.8 27.1 25.8 21.2
     $\Delta$ (HICRA - GRPO) -0.6 +0.3 +1.8 +2.1 -1.4 +0.9
Qwen2.5-7B-Base
@@
     ORZ (Hu et al., 2025) 18.8 14.8 80.2 52.4 39.7 45.9
     SimpleRL (Zeng et al., 2025) 16.7 3.3 78.2 42.8 34.9 38.2
     Entropy Regularization 16.0 9.3 57.4 50.3 33.1 40.6
@@
Table 2: Comparison of HICRA, GRPO on multimodal reasoning benchmarks.
VLM MathVista MathVision MathVerse EMMA
MiMO-VL-Instruct-2508
     Base 77.0 42.9 61.8 36.3
     GRPO 73.7 42.8 63.0 41.9
     HICRA 80.7 48.9 65.4 44.1
     $\Delta$ (HICRA - GRPO) +7.0 +6.1 +2.4 +2.2
Qwen2.5-VL-7B-Instruct
     Base 66.6 23.6 45.9 22.3
     GRPO 70.8 25.8 48.8 31.8
     HICRA 71.4 28.7 48.2 33.0
     $\Delta$ (HICRA - GRPO) +0.6 +2.9 -0.6 +1.2
Our primary results, summarized in Table 1, Table 2, show that HICRA consistently and outperforms both the GRPO baselines across text-only models and vision-language models on various benchmarks. On the strongest base model, Qwen3-4B-Instruct, HICRA’s gains demonstrate that even on highly capable models, selectively amplifying the learning signal for strategic reasoning yields substantial improvements. This trend holds for non-instruct-tuned models as well, providing strong empirical evidence for our central claim: by identifying and focusing on the emergent strategic bottleneck, HICRA accelerates the development of advanced reasoning abilities more efficiently than agnostic methods.
@@ 4.3.3 HICRA’s Limitation: Dependency on a Procedural Foundation
HICRA’s effectiveness is predicated on a key assumption: that the base model should readily possess a reasonable foundation for low-level procedural correctness. As shown in Figure 7, when this foundation is lacking – as observed with Llama-3.1-Instruct – HICRA can fail to provide an advantage over GRPO. Seen from the semantic entropy graph, there is a reverse trend between GRPO and HICRA, implying an opposite training focus on planning tokens and execution tokens. HICRA’s enforced strategic exploration becomes counterproductive if the model cannot reliably execute the plans it generates, leading to unstable learning dynamics and learning effects observed on Llama-3.1. This suggests that HICRA is most effective when applied to models that have already achieved a degree of procedural reliability, highlighting an important dependency for future work on more adaptive, model-aware hierarchical methods.
4.4 Semantic Entropy: A Compass for Strategic Exploration
@@
As shown in Figure 8 for MiMO-VL-7B, token-level entropy “collapses” for both HICRA and GRPO, simply because the vast majority of low-level tokens are doomed to become certain, thus pulling the average token entropy down. However, this decrease in token entropy might mislead researchers to suggest that exploration has ceased. Similarly, the Pass@8 (Training) metric quickly saturates, rendering it useless for distinguishing the ongoing learning dynamics.
@@
Recent work has proposed high-entropy tokens, sometimes called “fork tokens,” to imply its role as proxies for decision points in a reasoning trace (Wang et al., 2025d). Our analysis investigates the relationship between our functionally-defined planning tokens and this entropy-based definition.
@@
Figure 10: Planning Tokens, High-Entropy Tokens and Shared Tokens are highlighted with different colors. This concrete example suggests how these two definitions differ: Planning Tokens function as strategic skeletons of a reasoning solution and are thus sparse, with more than half of these semantic units also having higher entropy. In contrast, a majority of high-entropy tokens only exhibits high-variations in its phrasing, spreading across low-level executions and high-level planning. Fig. 9 reveals that less than 10% high-entropy tokens serve the semantic function of planning.
@@ 5 Related Works
Reinforcement Learning for LLM Reasoning. The application of Reinforcement Learning (RL) has been pivotal in enhancing the complex reasoning abilities of Large Language Models (LLMs). Seminal work by Ouyang et al. demonstrated the effectiveness of learning from human feedback to align models with user instructions. More recently, algorithms like Group Reward Policy Optimization (Guo et al., 2025) have been developed to specifically incentivize reasoning capabilities in LLMs,VLMs, Agents (Liu et al., 2025b; Yu et al., 2025; Team et al., 2025; Liu et al., 2025a; Wang et al., 2025c; b; Su et al., 2025; Dai et al., 2025; Zheng et al., 2025), leading to significant performance gains on downstream performance. While these methods have proven empirically successful, they typically apply optimization pressure agnostically across all generated tokens, without distinguishing between different functional roles within the reasoning process. Our work builds on this foundation but introduces a more targeted approach by focusing on the emergent reasoning hierarchy.
@@
Furthermore, recent work has identified high-entropy ”fork tokens” as potential proxies for critical decision points in reasoning (Wang et al., 2025d). Our work distinguishes itself by defining planning tokens based on their semantic function. We also validate the limitation of identifying crucial tokens solely based on entropy.
@@
Hierarchical Reasoning and Cognition. The concept of hierarchical processing is a cornerstone of cognitive neuroscience, which posits that the human brain separates high-level, abstract planning from low-level motor or procedural execution (Huntenburg et al., 2018; Murray et al., 2014; Zeraati et al., 2023; Zhu et al., 2025; Xiong et al., 2025). HRM (Wang et al., 2025a) is inspired this cognitive architecture to design a specific neural architecture for hierarchical reasoning. Concurrently, this cognitive model provides a compelling parallel to the functional hierarchy we identify in RL-tuned LLMs, proposing that LLMs similarly develop a functional separation between strategic planning and procedural execution.
@@ 6 Conclusion and Promising Directions
Our findings open several promising avenues for future research, suggesting a paradigm shift in how we approach RL for language models.
Rethinking Action in RL for LMs: Our analysis indicates that the token, as the fundamental unit of action, is too granular and often lacks semantic weight. A single token rarely determines the subsequent logical direction of a complex thought process. Future work should explore a higher-level action space, where an ”action” could be a semantic concept, a reasoning step (e.g., ’apply theorem X’), or a complete phrasal unit (n-gram). This would align the optimization process more closely with the cognitive steps of reasoning and could lead to more stable and efficient learning.
Dynamically Adaptive Hierarchical Methods: HICRA focused learning efforts on the strategic bottleneck. A more powerful approach should be capable of diagnosing the current learning bottleneck, which can adaptively shift optimization pressure between procedural execution and strategic planning based on real-time diagnostics of the model’s weaknesses, perhaps through a curriculum learning framework or meta-learning.
Finer-Grained, Process-Oriented Rewards: This work justifies a move away from sparse, outcome-based rewards (correct vs. incorrect final answer) towards process-oriented reward models. Such models could provide fine-grained credit assignment by, for example, rewarding a brilliant strategic choice even if it is marred by a minor calculation error. This decouples the evaluation of planning from execution, allowing the model to learn both skills more effectively.
@@
Generalizing the Reasoning Hierarchy Beyond Mathematics: The functional hierarchy between planning and execution is likely a universal feature of complex, multi-step generation. Investigating whether similar dynamics emerge in domains like code generation , agentic tool-use, long-form generation, etc, could yield novel, domain-specific optimization strategies.
Appendix A Full Training Dynamics
Figure 11: Training Dynamics across different LLMs and VLMs.
Following the discussion in the main paper, we make the following further observations based on the provided training charts:
  • For some models, the initial skill-consolidation phase is brief or absent. In the cases of the Vision-Language Models (Qwen2.5 VL-Instruct and MiMO VL-Instruct) and Qwen3 4B-Instruct, the exploration of strategic planning begins almost immediately at the start of training. This is evidenced by a significant and immediate rise in the semantic entropy of strategic grams, which occurs in tandem with a rapid boost in validation accuracy. We conjecture this is because: (a) for the VL scenarios, publicly available datasets is learned quickly by state-of-the-art models (Wang et al., 2025c); (b) strong base models like Qwen3 4B-Instruct already possess a solid foundation of low-level skills and primarily need to adapt to formatting before focusing on higher-level strategic planning.
  • Token-level entropy does not directly correlate with model accuracy. This is strongly supported across multiple experiments. For instance, with Llama3.1 8B, Qwen3 4B-Instruct, and the VL models, token-level entropy either remains flat or decreases throughout training. During the same period, however, validation accuracy shows a steady and significant increase. This demonstrates a clear disconnect between next-token uncertainty and overall task performance.
  • Token-level entropy is misleading for policy exploration. This observation holds true across all experiments. The Qwen3 4B-Instruct model offers a particularly stark example: its token-level entropy remains almost perfectly flat, while its semantic entropy (diversity of strategic grams) consistently increases throughout training. This contrast highlights that the variety of semantic structures a model learns is completely different from the statistical uncertainty of its next-token predictions.
    The core difference is about scale: token-level entropy measures the uncertainty of every next-token, including low-level tokens such as formatting, executions that are doomed to become confident throught training. In contrast, semantic entropy measures the diversity of the overall ideas being expressed. A model can be very predictable in its next-token choice under a given context but still create a wide variety of different arguments or structures.
Appendix B The Distribution Matching Perspective of Policy Gradients
Imagine an ideal, or ”target,” policy, $\pi^{*}(a|s)$, that we want our current policy, $\pi_{\theta}(a|s)$, to emulate. We can conceptualize this target distribution as being proportional to the exponentiated advantage of the actions:
$\pi^{*}(a|s)\propto\pi_{\theta_{old}}(a|s)\exp(\hat{A}(a,s))$ (1)
Or more concretely,
$\pi^{*}(a|s)=\frac{1}{Z(s)}\pi_{\theta_{old}}(a|s)\exp\left(\frac{\hat{A}(a,s)}{\beta}\right)$ (2)
This target policy, $\pi^{*}(a|s)$, re-weights the old policy based on the advantage of each action. Here, actions with a positive advantage ($\hat{A}>0$) get their probability boosted exponentially, while actions with a negative advantage ($\hat{A}<0$) get their probability suppressed. The term $\beta$ acts as a ”temperature” parameter.
The goal is to find a new policy, $\pi_{\theta}$, that is as close as possible to this ideal target distribution, $\pi^{*}$. This is equivalent to minimizing the KL divergence:
$\min_{\theta}\text{KL}(\pi^{*}(a|s)||\pi_{\theta}(a|s))$ (3)
Expanding this KL divergence term:
$\text{KL}(\pi_{\theta}||\pi^{*})$ $=\mathbb{E}_{a\sim\pi_{\theta}}\left[\log\frac{\pi_{\theta}(a|s)}{\pi^{*}(a|s)}\right]$
$=\mathbb{E}_{a\sim\pi_{\theta}}[\log\pi_{\theta}(a|s)-\log\pi^{*}(a|s)]$
Substitute our definition of $\log\pi^{*}(a|s)=\log\pi_{\theta_{old}}(a|s)+\frac{\hat{A}(a,s)}{\beta}-\log Z(s)$:
$=\mathbb{E}_{a\sim\pi_{\theta}}\left[\log\pi_{\theta}(a|s)-\left(\log\pi_{\theta_{old}}(a|s)+\frac{\hat{A}(a,s)}{\beta}-\log Z(s)\right)\right]$
$\propto\mathbb{E}_{a\sim\pi_{\theta}}\left[\left(\log\pi_{\theta}(a|s)-\log\pi_{\theta_{old}}(a|s)\right)-\frac{\hat{A}(a,s)}{\beta}\right]$
$=\frac{1}{\beta}\mathbb{E}_{a\sim\pi_{\theta}}\left[\beta\text{KL}(\pi_{\theta}||\pi_{\theta_{old}})-\hat{A}(a,s)\right]$
Minimizing this is equivalent to maximizing its negative:
$\max_{\theta}\mathbb{E}_{a\sim\pi_{\theta}}\left[\hat{A}(a,s)-\beta\text{KL}(\pi_{\theta}||\pi_{\theta_{old}})\right]$ (4)
This expression is nearly identical to the PPO-KL objective, where the policy update is constrained using a KL divergence regularizer (Schulman et al., 2017).
Therefore, the PPO objective is essentially solving a distribution matching problem toward a target distribution shaped by the advantage function.